OpenStudy (anonymous):

Find the derivative of h(x)=(x²-x-12)/(x²+25) So I know the functions factor out to be (x+3)(x−4)/(x²+25) So far I have this [(x²+25)(2x - 1)]-[(x²-x-12)(2x )] / (x²+25)² I seriously don't even know if this is right? What steps proceed after?? I have no clue I tried to figure it out and it went completely wrong! I am so stuck on the derivatives. This is my last resort, can anyone show me how this should be answered correctly? My brain is fried...

OpenStudy (anonymous):

from what i have learned you need to bring that bottom term to the numerator.

OpenStudy (anonymous):

d(u/v)/dx=(u'v-v'u)/v² u'v=(2x-1)(x²+25) vu'=2x(x²-x-12) v²=x^4+50x²+625 Simplify as required

OpenStudy (cggurumanjunath):

what you have done is perfectly alright @Ikushinkan .